OpenStudy (anonymous):

Which equation is true for the value b = 2?

OpenStudy (antiprime):

ok so u want a equation that makes this true

OpenStudy (antiprime):

what are the choices

OpenStudy (antiprime):

@summer335

OpenStudy (antiprime):

b-2=0 lol

OpenStudy (antiprime):

2b-4=0 3b-6=0 4b-8=0 5b-10=0 6b-12=0

OpenStudy (antiprime):

i need the choices

OpenStudy (antiprime):

@summer335 i cant do this with out knowing what i have to chooose from